Find the values of Cp and ᵞ for a gas whose gas




constant is 0.324 kJ/kg K and Cv = 0 84 kJ/kg K.




If 2.25 kg of this gas with an initial volume of




1.15 m3 undergoes a process during which its




pressure remains constant at 7 bar and its




temperature becomes 280‘C at the end of the




process, calculate: (a) the change in internal




energy, (b) the change in enthalpy, (c) the




transferred heat, and (d) the work done.

1
Expert's answer
2021-12-13T11:05:56-0500

Internal energy


"U=mC_v∆T 
\\U=2.25\times0.84\times553=1045.17KJ"

Enthalpy

"1bar=10^5 pa"


"H=U+PV 
\\H=1045.17+7\times10^5\times1.15 
\\H=1045.17+805\\H=1850.17KJ"

Work done (W)=PdV

W="7\times10^5\times1.15=805KJ"